1. What is the first step to solving the equation 1= square root of x-3
1
Expert's answer
2021-03-16T07:07:57-0400

The first step to solve the equation,

1 = "\sqrt{x-3}" ,

is to square both the sides.

i.e., 1"^2" = "(\sqrt{x - 3})^2"

1 = x - 3

x = 1 + 3

"\therefore" x = 4